PLEASE HELP ME SOLVE THIS WORD PROBLEM: "AGE"
A FATHER IS 30 YEARS OLDER THAN HIS SON. IN 12 YEARS, THE FATHER WILL BE TWICE AS OLD AS HIS SON THEN. FIND THEIR PRESENT AGES.

Answer by TimothyLamb(4379)

f = s + 30
f + 12 = 2(s + 12)
---
f + 12 = 2s + 24
s + 30 + 12 = 2s + 24
s = 18
Answer:
f = 48
s = 18
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php